# Tomographic Alcock-Paczynski Test with Marked Correlation Functions

### v4

@@ I Introduction
Over the past two decades, the Stage-III galaxy surveys, such as [2df:Colless:2003wz, beutler20116df, blake2011wigglez, blake2011wigglezb, york2000sloan, Eisenstein:2005su, Percival:2007yw, anderson2012clustering, alam2017clustering], have expanded the boundaries of our understanding of the universe. Ongoing and upcoming in the next decade, advanced optical telescopes, such as the Dark Energy Spectroscopic Instrument (DESI) [^1] https://desi.lbl.gov/, the Vera C. Rubin Observatory (LSST) [^2] https://www.lsst.org/, the Euclid satellite [^3] http://sci.esa.int/euclid/, the Roman Space Telescope [^4] https://roman.gsfc.nasa.gov/, and the Chinese Space Station Telescope (CSST) [^5] http://nao.cas.cn/csst/, will conduct Stage-IV surveys. These surveys will accurately map the distribution of cosmic structures across a spatial scale of 10 billion light-years and characterize cosmic evolution over 10 billion years. Through statistical analysis of the upcoming survey data, we aim to understand the nature of dark matter and dark energy with a precision higher than 1%. In particular, the precision of these surveys will allow us to explore non-linear clustering regions of galaxies, which contain a wealth of information. This poses a significant challenge for the cosmological analysis of the data. We need new statistical analysis methods that can effectively handle the nonlinear region, in order to comprehensively utilize the observational results and obtain more physical information.
The Alcock-Paczynski (AP) test [AP1979] serves as a means to probe the expansion history of the Universe. In essence, the test can be applied to any structure in the Universe displaying random orientations, such as halos, filaments, or walls, leading to a spherically symmetric ensemble average. Under a specific cosmological model, the radial and tangential sizes of distant objects or structures can be expressed as $\Delta r_{\parallel}$ and $\Delta r_{\perp}$ (Eq.2). If the models deviate from the true cosmology, the computation of $D_{A}$ and $H(z)$ results in an incorrect estimation of $\Delta r_{\parallel}$ and $\Delta r_{\perp}$. This inaccuracy manifests as geometric distortions along the line of sight (LOS) and perpendicular directions, referred to as the AP distortions. The AP distortions can be measured and quantified through statistical analysis of the large-scale galaxy distribution. Consequently, the observed distortions offer a method to constrain cosmological parameters.
@@
A key limitation of the current tomographic AP method is its exclusive reliance on two-point statistics. While this approach is well-established, two-point statistics may be insufficient for extracting cosmological information from non-Gaussian fields. This limitation is particularly relevant for upcoming stage-IV surveys, which will probe the non-linear clustering regime where the density field becomes highly non-Gaussian. To enhance its constraining power, a logical extension would be to apply the AP method to statistics beyond 2PCF and power spectrum. A recent addition to this field is the density-marked correlation functions (MCFs)[stoyan_correlations_1984], offering a density-dependent statistic that extends beyond two-point statistics and can effectively capture non-Gaussian features [Beisbart2000, Beisbart2002, Gottl2002, Sheth:2004vb, Sheth:2005aj, Skibba2006, White_2009, White_2016, Satpathy:2019nvo, 2020arXiv200111024M, PMS2020, Massara:2022zrf, Jung:2024esv, Philcox:2020fqx, Philcox:2020srd, Ebina:2024zkv, Massara:2024cvu]. The MCFs introduce a weight for each galaxy using a “marker” based on its local density, thereby providing density-dependent clustering information. The conventional 2PCF approach is, of course, a specific case of MCFs, where all galaxies receive identical weights irrespective of differences in their physical properties and environments. The distinctive feature of MCFs lies in its ability to highlight statistics in regions of higher or lower density. We have the freedom to choose weights that allow the statistic to respond to both positive and negative powers of the density, making MCFs highly versatile. More recently, studies using N-body simulations have demonstrated that combining different weights in MCFs can significantly improve cosmological parameter constraints [Yang:2020ysv, XiaoYuan_2022]. Based on these findings, [Lai:2023dzp, Yin:2024iqp] applied MCFs to SDSS galaxy data, further confirming enhanced parameter constraints compared to traditional methods.
@@
Another growing challenge in modern surveys is the increasingly complexity of covariance matrix estimation. This is mainly due to two aspects: 1) as surveys advance we must use more complex statistics, also for the wider redshift coverage of the stage-IV surveys we need more redshift bins, this both increases data vector sizes and makes covariance calculations much harder; 2) larger volume and higher precision survey samples require larger and higher resolution samples for covariance matrix, making simulation samples preparation particularly difficult and expensive. In this study, the combined MCFs also leads to significant expansion of size of statistics. To solve this problem, we adopt a compression scheme based on principal component analysis (PCA) [Huterer_2003, Crittenden_2009] to address this issue. PCA is an efficient technique to reduce the size of statistics of combining different types of MCFs in this study. Besides, using PCA we can also compress the full $(s,\mu)$-dependence of the correlation functions, which is extremely difficult to utilize in the previous cosmological studies work due to the extremely difficult covariance estimation.
@@
The redshift uncertainties in slitless spectroscopy may also impact the performance of the tomographic AP test. Tomographic AP test is sensitive to the intrinsic anisotropy in the survey data. In the next few years, slitless instruments will be used in CSST [Gong_2019] and the Euclid satellite [EUCLID] to determine the spectroscopic redshifts of galaxies. While slitless spectroscopy enhances the efficiency and practicality of spectroscopic surveys, it introduces significantly larger redshift errors compared to conventional fiber-based methods. Our analysis in  [Xiao_2022] shows the tomographic AP test’s constraining power experiences mild degradation under conservative CSST assumptions ($\sigma_{z}\lesssim 0.002$). In this study, we further test the impact of redshift errors caused by slitless instruments in the cases of MCF, and the cases of MCFs and ($s,\mu$)-dependency 2-dimensional MCFs with PCA compression. Following [Xiao_2022], in this study, we will not only test the impact of the conservative estimate for CSST, but also test a larger estimate and an estimate with a deviation of linear between errors and redshifts.
@@
In this study, utilizing a large N-body simulation, we will, for the first time, assess the performance of MCFs in constraining cosmological parameters, with a focus on $\Omega_{m}$ and $w$, using the tomographic AP test. The paper is structured as follows: In Sect. II, we introduce the simulation data used for our analysis. Sect. III presents the methodology for calculating the MCF on mock data and the tomographic AP test. In Sect. IV, we present our findings and compare the constraints between the standard 2PCF and MCFs. We also highlight constraints improved by using the combined MCFs with different weights and the PCA compression method. As one of the follow-up works of [Xiao_2022], the impact of redshift error from slitless instruments on combined MCFs and PCA compression is also considered. Besides, a robustness check about the systematics estimation is attached to clarify the current limitations of our method. Finally, we summarize our results and conclude in Sect. V.
@@ II Data
Throughout this analysis, we use the halos and subhalos identified by the ROCKSTAR halo finder [ROCKSTAR] at snapshots $z_{1}=0.607$ and $z_{2}=1$, resulting in halo catalogs comprising snapshots generated from the simulated particles. It is important to note that when dealing with real observational data, it becomes essential to incorporate a lightcone approach to model systematics effectively [LI16]. However, in this study, where our primary concern is the impact of redshift errors arising from RSD, we opt for simplicity and utilize snapshots.
@@ III.1 Tomographic AP test
Ref. [LI14] introduced a tomographic analysis of small-scale galaxy clustering aimed at effectively distinguishing the AP effect from RSD. They employed the integrated 2-dimensional 2PCF along various LOS to evaluate the level of anisotropy across different redshifts. This is represented by
@@
We adopt the widely used Landy–Szalay estimator [Landy] to calculate the 2PCF
@@
For real observations, it is beneficial to adopt multiple pairs of redshift slices to enhance constraining power. While various redshift slicing approaches are feasible [LI16, LI19], the specific selection employed in this study, with $z_{1}=1.0$ and $z_{2}=0.6069$, is indicative of stage-IV surveys that typically obtain numerous spectroscopic galaxy samples around $z\sim 1$. For CSST, the slitless redshift errors may become large at $z\gtrsim 1$ due to its potential inability to detect Ly-$\alpha$ lines in galaxies with redshifts greater than 1. Additionally, [Ma_2020] demonstrated that the systematics of the tomographic AP method are notably larger for redshift slices $z\gtrsim 1$. Therefore, we opt not to include additional tomography slices at higher redshifts.
Additionally, effects like the RSD contribute to a non-zero redshift evolution of anisotropy in $\delta\widehat{\xi}_{\Delta s}$. Furthermore, apart from the RSD effect, various sources of systematics originate from observational effects like selection bias and fiber collision. To avoid the complexities associated with modeling these systematics, a practical approach involves correcting for them using mock surveys [LI14, LI15, LI16]. In this context, $\delta\widehat{\xi}_{\Delta s}^{\rm corr}(\mu)$ represents the redshift dependence after correcting for systematics for the selected two redshift bins, defined as
@@
Following [Xiao_2022], the covariance matrix $\boldsymbol{C}$ is estimated from the $\mathtt{BIGMD}$ simulation. We initially divide the entire simulation box into $8^{3}$ sub-boxes, each with a side length of 325 $h^{-1}{\rm Mpc}$. Subsequently, we employ empirical covariance matrix estimation based on these 512 sub-boxes. The number of sub-box samples is larger than the size of the $\boldsymbol{p}$, which ensures the convergence of the covariance matirx. We do a robustness check of covariance matrix in Appendix B. The statistical uncertainty of the correction term is also negligible compared to $\delta\widehat{\xi}_{\Delta s}(z_{1},z_{2},\mu)$ (see Eq. 8). Therefore, the covariance only accounts for the contribution from the fluctuations of $\delta\widehat{\xi}_{\Delta s}(z_{1},z_{2},\mu)$. Note that we choose sub-boxes located at different spatial positions to compute $\delta\widehat{\xi}_{\Delta s}(z_{1},z_{2},\mu)$. In this case, the comparison between two redshfit bins is always made for sub-samples with different cosmic variance, so that to avoid the effect of related noise fluctuations and structural features between the two snapshots.
@@ III.2 MCF statistics
MCFs extend the standard 2PCF, a statistical measure assigning weights to objects before measuring their correlation [White_2016, Yang:2020ysv, XiaoYuan_2022]. Physically, we assign weights based on their local densities to consider environmental dependence in our analysis. MCFs are sensitive to the chosen weight, as clustering and redshift space distortion effects exhibit notable variations in dense and sparse regions of the universe. As outlined in [Yang:2020ysv], these weights are designed in the following way
@@
where $\rho_{n_{\rm NB}}(\boldsymbol{x})$ is the environmental number density around a specific galaxy located at $\boldsymbol{x}$, and $W_{k}$ is the smoothing kernel, for which we choose the third-order B-spline functions which has non-zero values within a sphere of radius $h_{W}$ $h^{-1}{\rm Mpc}$. In this study, we set $n_{\rm NB}=30$, indicating that the radius $h_{w}$ is determined by the distance between a selected halo and its 30-th nearest neighbor [Gingold1977, Lucy1977].
@@
The 2PCF and MCFs of the halo catalogs within backgrounds of wrong cosmologies are obtained by applying the mapping approach using coordinate transformations in the level of 2PCF and MCFs instead of mock data, as detailed in Appendix A(especially, Eq. A1 and Fig. 6) of [LI18]:
@@
where $\kappa_{\|}\equiv D_{A,\rm target}/D_{A,\rm fid}$ and $\kappa_{\perp}=H_{\rm fid}/H_{\rm target}$ and they are computed in the two snapshots of the target and fiducial cosmologies, fixing $\mathrm{Mpc}/h$ unit, respectively. This approach is much more efficient than converting the halo samples into different backgrounds and remeasuring the 2PCFs and MCFs. However, it should be noted that: i) changes in the values of the weights are not captured by Eq. 20. This discrepancy arises because the AP effect nonuniformly distorts the geometry, causing the set of $n_{\rm NB}$ nearest neighbors to differ from one cosmology to another; ii) only the background evolution of different cosmologies is considered, ignoring the effects arising from structure growth. However, the former has been shown in previous studies [Yang:2020ysv] to have only a very minor impact, which can be safely ignored. The latter is not to be considered in the tomographic AP test, as the AP effect is the geometric distortion corresponding to the background evolution itself and we only utilize the background evolution of different cosmologies to constrain cosmology.
When constraining cosmological parameters, [Yang:2020ysv] has been observed that the constraining power of MCFs combining multiple different $\alpha$ values is significantly stronger than the results obtained using a single $\alpha$. Due to this observation, we will explore the combined MCFs with different $\alpha$, forming $\boldsymbol{p}^{T}=[\boldsymbol{p}_{\alpha_{1}}^{T},\boldsymbol{p}_{\alpha_{2}}^{T},\cdots,\boldsymbol{p}_{\alpha_{n}}^{T}]$, with check the ordering in covariance
@@
In this work, the full 2-dimensional 2PCF and MCFs are also applied to constrain cosmology parameters with PCA (see Sec.III.3 for details). The full 2-dimensional 2PCF is defined as [Park:2019mvn]

### v5

@@ I Introduction
Over the past two decades, the Stage-III galaxy surveys, such as [13, 8, 9, 10, 68, 19, 47, 4, 2], have expanded the boundaries of our understanding of the universe. Ongoing and upcoming in the next decade, advanced optical telescopes, such as the Dark Energy Spectroscopic Instrument (DESI) [^1] https://desi.lbl.gov/, the Vera C. Rubin Observatory (LSST) [^2] https://www.lsst.org/, the Euclid satellite [^3] http://sci.esa.int/euclid/, the Roman Space Telescope [^4] https://roman.gsfc.nasa.gov/, and the Chinese Space Station Telescope (CSST) [^5] http://nao.cas.cn/csst/, will conduct Stage-IV surveys. These surveys will accurately map the distribution of cosmic structures across a spatial scale of 10 billion light-years and characterize cosmic evolution over 10 billion years. Through statistical analysis of the upcoming survey data, we aim to understand the nature of dark matter and dark energy with a precision higher than 1%. In particular, the precision of these surveys will allow us to explore non-linear clustering regions of galaxies, which contain a wealth of information. This poses a significant challenge for the cosmological analysis of the data. We need new statistical analysis methods that can effectively handle the nonlinear region, in order to comprehensively utilize the observational results and obtain more physical information.
The Alcock-Paczynski (AP) test [3] serves as a means to probe the expansion history of the Universe. In essence, the test can be applied to any structure in the Universe displaying random orientations, such as halos, filaments, or walls, leading to a spherically symmetric ensemble average. Under a specific cosmological model, the radial and tangential sizes of distant objects or structures can be expressed as $\Delta r_{\parallel}$ and $\Delta r_{\perp}$ (Eq.2). If the models deviate from the true cosmology, the computation of $D_{A}$ and $H(z)$ results in an incorrect estimation of $\Delta r_{\parallel}$ and $\Delta r_{\perp}$. This inaccuracy manifests as geometric distortions along the line of sight (LOS) and perpendicular directions, referred to as the AP distortions. The AP distortions can be measured and quantified through statistical analysis of the large-scale galaxy distribution. Consequently, the observed distortions offer a method to constrain cosmological parameters.
@@
A key limitation of the current tomographic AP method is its exclusive reliance on two-point statistics. While this approach is well-established, two-point statistics may be insufficient for extracting cosmological information from non-Gaussian fields. This limitation is particularly relevant for upcoming stage-IV surveys, which will probe the non-linear clustering regime where the density field becomes highly non-Gaussian. To enhance its constraining power, a logical extension would be to apply the AP method to statistics beyond 2PCF and power spectrum. A recent addition to this field is the density-marked correlation functions (MCFs)[57], offering a density-dependent statistic that extends beyond two-point statistics and can effectively capture non-Gaussian features [7, 6, 23, 55, 54, 56, 61, 62, 53, 43, 50, 42, 27, 49, 48, 18, 41]. The MCFs introduce a weight for each galaxy using a “marker” based on its local density, thereby providing density-dependent clustering information. The conventional 2PCF approach is, of course, a specific case of MCFs, where all galaxies receive identical weights irrespective of differences in their physical properties and environments. The distinctive feature of MCFs lies in its ability to highlight statistics in regions of higher or lower density. We have the freedom to choose weights that allow the statistic to respond to both positive and negative powers of the density, making MCFs highly versatile. More recently, studies using N-body simulations have demonstrated that combining different weights in MCFs can significantly improve cosmological parameter constraints [66, 65]. Based on these findings, [29, 67] applied MCFs to SDSS galaxy data, further confirming enhanced parameter constraints compared to traditional methods.
@@
Another growing challenge in modern surveys is the increasingly complexity of covariance matrix estimation. This is mainly due to two aspects: 1) as surveys advance we must use more complex statistics, also for the wider redshift coverage of the stage-IV surveys we need more redshift bins, this both increases data vector sizes and makes covariance calculations much harder; 2) larger volume and higher precision survey samples require larger and higher resolution samples for covariance matrix, making simulation samples preparation particularly difficult and expensive. In this study, the combined MCFs also leads to significant expansion of size of statistics. To solve this problem, we adopt a compression scheme based on principal component analysis (PCA) [26, 14] to address this issue. PCA is an efficient technique to reduce the size of statistics of combining different types of MCFs in this study. Besides, using PCA we can also compress the full $(s,\mu)$-dependence of the correlation functions, which is extremely difficult to utilize in the previous cosmological studies work due to the extremely difficult covariance estimation.
@@
The redshift uncertainties in slitless spectroscopy may also impact the performance of the tomographic AP test. Tomographic AP test is sensitive to the intrinsic anisotropy in the survey data. In the next few years, slitless instruments will be used in CSST [22] and the Euclid satellite [31] to determine the spectroscopic redshifts of galaxies. While slitless spectroscopy enhances the efficiency and practicality of spectroscopic surveys, it introduces significantly larger redshift errors compared to conventional fiber-based methods. Our analysis in  [64] shows the tomographic AP test’s constraining power experiences mild degradation under conservative CSST assumptions ($\sigma_{z}\lesssim 0.002$). In this study, we further test the impact of redshift errors caused by slitless instruments in the cases of MCF, and the cases of MCFs and ($s,\mu$)-dependency 2-dimensional MCFs with PCA compression. Following [64], in this study, we will not only test the impact of the conservative estimate for CSST, but also test a larger estimate and an estimate with a deviation of linear between errors and redshifts.
@@
In this study, utilizing a large N-body simulation, we will, for the first time, assess the performance of MCFs in constraining cosmological parameters, with a focus on $\Omega_{m}$ and $w$, using the tomographic AP test. The paper is structured as follows: In Sect. II, we introduce the simulation data used for our analysis. Sect. III presents the methodology for calculating the MCF on mock data and the tomographic AP test. In Sect. IV, we present our findings and compare the constraints between the standard 2PCF and MCFs. We also highlight constraints improved by using the combined MCFs with different weights and the PCA compression method. As one of the follow-up works of [64], the impact of redshift error from slitless instruments on combined MCFs and PCA compression is also considered. Besides, a robustness check about the systematics estimation is attached to clarify the current limitations of our method. Finally, we summarize our results and conclude in Sect. V.
@@ II Data
Throughout this analysis, we use the halos and subhalos identified by the ROCKSTAR halo finder [5] at snapshots $z_{1}=0.607$ and $z_{2}=1$, resulting in halo catalogs comprising snapshots generated from the simulated particles. It is important to note that when dealing with real observational data, it becomes essential to incorporate a lightcone approach to model systematics effectively [35]. However, in this study, where our primary concern is the impact of redshift errors arising from RSD, we opt for simplicity and utilize snapshots.
@@ III.1 Tomographic AP test
Ref. [34] introduced a tomographic analysis of small-scale galaxy clustering aimed at effectively distinguishing the AP effect from RSD. They employed the integrated 2-dimensional 2PCF along various LOS to evaluate the level of anisotropy across different redshifts. This is represented by
@@
We adopt the widely used Landy–Szalay estimator [30] to calculate the 2PCF
@@
For real observations, it is beneficial to adopt multiple pairs of redshift slices to enhance constraining power. While various redshift slicing approaches are feasible [35, 33], the specific selection employed in this study, with $z_{1}=1.0$ and $z_{2}=0.6069$, is indicative of stage-IV surveys that typically obtain numerous spectroscopic galaxy samples around $z\sim 1$. For CSST, the slitless redshift errors may become large at $z\gtrsim 1$ due to its potential inability to detect Ly-$\alpha$ lines in galaxies with redshifts greater than 1. Additionally, [40] demonstrated that the systematics of the tomographic AP method are notably larger for redshift slices $z\gtrsim 1$. Therefore, we opt not to include additional tomography slices at higher redshifts.
Additionally, effects like the RSD contribute to a non-zero redshift evolution of anisotropy in $\delta\widehat{\xi}_{\Delta s}$. Furthermore, apart from the RSD effect, various sources of systematics originate from observational effects like selection bias and fiber collision. To avoid the complexities associated with modeling these systematics, a practical approach involves correcting for them using mock surveys [34, 32, 35]. In this context, $\delta\widehat{\xi}_{\Delta s}^{\rm corr}(\mu)$ represents the redshift dependence after correcting for systematics for the selected two redshift bins, defined as
@@
Following [64], the covariance matrix $\boldsymbol{C}$ is estimated from the $\mathtt{BIGMD}$ simulation. We initially divide the entire simulation box into $8^{3}$ sub-boxes, each with a side length of 325 $h^{-1}{\rm Mpc}$. Subsequently, we employ empirical covariance matrix estimation based on these 512 sub-boxes. The number of sub-box samples is larger than the size of the $\boldsymbol{p}$, which ensures the convergence of the covariance matirx. We do a robustness check of covariance matrix in Appendix B. The statistical uncertainty of the correction term is also negligible compared to $\delta\widehat{\xi}_{\Delta s}(z_{1},z_{2},\mu)$ (see Eq. 8). Therefore, the covariance only accounts for the contribution from the fluctuations of $\delta\widehat{\xi}_{\Delta s}(z_{1},z_{2},\mu)$. Note that we choose sub-boxes located at different spatial positions to compute $\delta\widehat{\xi}_{\Delta s}(z_{1},z_{2},\mu)$. In this case, the comparison between two redshfit bins is always made for sub-samples with different cosmic variance, so that to avoid the effect of related noise fluctuations and structural features between the two snapshots.
@@ III.2 MCF statistics
MCFs extend the standard 2PCF, a statistical measure assigning weights to objects before measuring their correlation [62, 66, 65]. Physically, we assign weights based on their local densities to consider environmental dependence in our analysis. MCFs are sensitive to the chosen weight, as clustering and redshift space distortion effects exhibit notable variations in dense and sparse regions of the universe. As outlined in [66], these weights are designed in the following way
@@
where $\rho_{n_{\rm NB}}(\boldsymbol{x})$ is the environmental number density around a specific galaxy located at $\boldsymbol{x}$, and $W_{k}$ is the smoothing kernel, for which we choose the third-order B-spline functions which has non-zero values within a sphere of radius $h_{W}$ $h^{-1}{\rm Mpc}$. In this study, we set $n_{\rm NB}=30$, indicating that the radius $h_{w}$ is determined by the distance between a selected halo and its 30-th nearest neighbor [21, 38].
@@
The 2PCF and MCFs of the halo catalogs within backgrounds of wrong cosmologies are obtained by applying the mapping approach using coordinate transformations in the level of 2PCF and MCFs instead of mock data, as detailed in Appendix A(especially, Eq. A1 and Fig. 6) of [36]:
@@
where $\kappa_{\|}\equiv D_{A,\rm target}/D_{A,\rm fid}$ and $\kappa_{\perp}=H_{\rm fid}/H_{\rm target}$ and they are computed in the two snapshots of the target and fiducial cosmologies, fixing $\mathrm{Mpc}/h$ unit, respectively. This approach is much more efficient than converting the halo samples into different backgrounds and remeasuring the 2PCFs and MCFs. However, it should be noted that: i) changes in the values of the weights are not captured by Eq. 20. This discrepancy arises because the AP effect nonuniformly distorts the geometry, causing the set of $n_{\rm NB}$ nearest neighbors to differ from one cosmology to another; ii) only the background evolution of different cosmologies is considered, ignoring the effects arising from structure growth. However, the former has been shown in previous studies [66] to have only a very minor impact, which can be safely ignored. The latter is not to be considered in the tomographic AP test, as the AP effect is the geometric distortion corresponding to the background evolution itself and we only utilize the background evolution of different cosmologies to constrain cosmology.
When constraining cosmological parameters, [66] has been observed that the constraining power of MCFs combining multiple different $\alpha$ values is significantly stronger than the results obtained using a single $\alpha$. Due to this observation, we will explore the combined MCFs with different $\alpha$, forming $\boldsymbol{p}^{T}=[\boldsymbol{p}_{\alpha_{1}}^{T},\boldsymbol{p}_{\alpha_{2}}^{T},\cdots,\boldsymbol{p}_{\alpha_{n}}^{T}]$, with check the ordering in covariance
@@
In this work, the full 2-dimensional 2PCF and MCFs are also applied to constrain cosmology parameters with PCA (see Sec.III.3 for details). The full 2-dimensional 2PCF is defined as [45]